OpenStudy (anonymous):

Ok, I need help with this: Can the statement be written as a true biconditional statement? If x = 5, then x^2 =25. What is a biconditional statement? I don't get it.

jimthompson5910 (jim_thompson5910):

A biconditional statement is something of the form If P, then Q AND if Q then P So this is essentially saying that P and Q are the same thing So are the statements x = 5 and x^2=25 the same statement?

OpenStudy (anonymous):

Yes, they are in the same statement. And....yeah....they are.

jimthompson5910 (jim_thompson5910):

not quite, they look the same, but we need to really check

jimthompson5910 (jim_thompson5910):

let's start with the first part: If P then Q So start with x = 5 and square both sides to get x^2 = 5^2 ---> x^2 = 25 So we've shown that if x = 5, then x^2 = 25. So that part works ------------- Now let's check: If Q then P So start with x^2 = 25 and take the square root of both sides to get x = 5 or x = -5 (remember the plus/minus when it comes to solving quadratics). But now we have the extra statement of x = -5, which is nowhere in the first statement. So this means that x^2 = 25 does NOT imply x = 5 alone So this part does not work Therefore, the two statements are NOT equivalent

OpenStudy (anonymous):

x=5 x^2=25 5^2=5*5 5*5=25 Isn't it right? Or am I missing something?

jimthompson5910 (jim_thompson5910):

If however, you wrote the first statement as | x | = 5, then the two would be equivalent Notice the difference now is that the x is surrounded by absolute value bars

jimthompson5910 (jim_thompson5910):

Yes but (-5)^2 = (-5)*(-5) = 25 as well

OpenStudy (anonymous):

Ok, thank you jim :)